OpenStudy (anonymous):

True/False: If \[3^{2x-1}=1 then 2x-1=log_3(2)\]

OpenStudy (anonymous):

False. It would be true if it were: \[3^{2x - 1} = 1\]then\[2x - 1 = \log_{3}(1)\] Remember the forms: \[y = b^{x}\]\[x = \log_{b}(y) \]

OpenStudy (anonymous):

okay, thank you :-)

OpenStudy (anonymous):

No Problem :)

OpenStudy (anonymous):

Oh, yikes... that suppose to be equal to 2 so is it true...

OpenStudy (anonymous):

Wait what? I don't understand.

OpenStudy (anonymous):

You mean it should be \[3^{2x - 1} = 2?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then it should be true.

OpenStudy (anonymous):

Okay, thanks!